OpenStudy (anonymous):

FAN AND MEDAL!! she drew a line of best fit passing through points (1,1) and (5,7) on the graph to show one way calculating the mean growth rate of the plants what is the slope of the line she drew?

OpenStudy (anonymous):

m=3/2

OpenStudy (anonymous):

Thanks @kvboogie

OpenStudy (anonymous):

Slope = (y2-y1)/(x2-x1) Slope = (7-1)/(5-1) = 6/4 = 3/2. 3/2 No Problem :)

OpenStudy (anonymous):

Fan ?

OpenStudy (anonymous):

sorry I thought I did?? @kvboogie